# DGLAP evolution at N3LO with the Candia algorithm

### v2

@@ I Introduction
The collinear substructure of the nucleon is described by parton distribution functions (PDFs), which map out the longitudinal momentum fractions of its inner constituent quarks and gluons in high-energy collisions. PDFs are essential components of factorization formulae in Quantum Chromodynamics (QCD), which allow for the perturbative calculation of hadronic cross sections. A precise and accurate determination of PDFs is a major endeavor [2, 3, 4] and is necessary to advance the current precision frontier of cross-section calculations at hadron colliders. In the QCD framework of collinear factorization, PDFs depend on two variables: the parton longitudinal momentum fraction $x$ and the energy scale $Q$ of the reaction. The PDF $x$-dependence cannot be predicted by first principles, at least within our current knowledge, and has to be extracted from global QCD analyses of experimental data for a wide variety of hadronic cross sections [2, 3, 4, 5, 6, 7]. The energy behavior of the PDFs instead, is predicted by perturbative QCD (pQCD) and is described by the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) [8, 9, 10, 11, 12] integro-differential renormalization group equations. The latter require knowledge of the $x$-dependent PDF functional form extracted from experimental data at an initial energy scale $Q_{0}$.
@@
Parton evolution is crucial for the success of all these programs and past and recent literature contains multiple efforts devoted to the development of algorithms using different approaches to solve DGLAP equations [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102]. In the past three decades, part of these methods have been implemented in efficient computer codes, some of which are publicly available. Examples of these codes are partonevolution [97] (Mellin-transform method with an optimized contour), Pegasus [103] ($U$-matrix method in Mellin space), Candia [104, 105] (logarithmic recursion relations in the $x$-space), Hoppet [106, 107] (Runge-Kutta method in the $x$-space), QCDNum [108] (polynomial spline interpolation in the $x$-space), and more recently Apfel [109, 110] (Runge-Kutta method and higher-order interpolations in the $x$-space), ChiliPDF [111] (global interpolation with Chebyshev polynomials in Mellin space), EKO [112] (evolution kernel operator method in Mellin space), uPDFevolv [113, 114] (parton branching method). A recent benchmarking exercise of QCD evolution at aN3LO is reported in Ref. [115].
@@ III N3LO Non-Singlet Exact Solutions in Power Series
$Z_{22}(x)$ $=\frac{1}{\beta_{3}\gamma}\left[32\pi^{2}[P^{(1)}(x)+16\pi r_{1}P^{(2)}(x)-8(\bar{c}+\bar{b}r_{1})P^{(3)}(x)\right],$
@@
The strategy we adopt when more than one recursion relation in Eqns. (37a)–(37d) is eligible for the computation of a specific $D^{s}_{t,m,n}$ coefficient, is based on following procedure:
@@ IV Truncated solutions for the DGLAP Singlet and Non-Singlet sectors at all orders in QCD
These recursion relations generate truncated solutions at arbitrary order $\kappa$ that are equivalent those obtained using the $U$-matrix method [103, 98].
@@ V Factorization/Renormalization Scale Dependence
$P^{(1)}\left(x,\frac{\mu_{F}^{2}}{\mu_{R}^{2}}\right)=$ $P^{(1)}(x)-\frac{\beta_{0}}{2}P^{(0)}(x)L$ (53)
$P^{(2)}\left(x,\frac{\mu_{F}^{2}}{\mu_{R}^{2}}\right)=$ $P^{(2)}(x)-\left[\beta_{1}P^{(0)}(x)+2\beta_{0}P^{(1)}(x)\right]L+\beta_{0}^{2}P^{(0)}(x)L^{2}$ (54)
$P^{(3)}\left(x,\frac{\mu_{F}^{2}}{\mu_{R}^{2}}\right)=$ $P^{(3)}-\left[\beta_{2}P^{(0)}+2\beta_{1}P^{(1)}(x)+3\beta_{0}P^{(2)}(x)\right]L$
$+\left[\frac{5}{2}\beta_{0}\beta_{1}P^{(0)}(x)+3\beta_{0}^{2}P^{(1)}(x)\right]L^{2}-\beta_{0}^{3}P^{(0)}(x)L^{3}\,$, (55)
@@ VI Heavy-Quark matching conditions
$q_{h}^{(n_{f}+1)}(x)=\bar{q}^{(n_{f}+1)}(x)=\frac{1}{2}\Big\{[A_{hq}^{\mathrm{S}}\otimes q^{+,(n_{f})}](x)+[A_{hg}^{\mathrm{S}}\otimes g^{(n_{f})}](x)\Big\}\,$, (61)
where we used the relation $q_{h}^{(n_{f}+1)}(m_{h})=\bar{q}_{h}^{(n_{f}+1)}(m_{h})$. The OMEs have the perturbative expansions given below
@@
$A^{\textrm{PS}}_{qq,h}(x)$ $=\left(\frac{\alpha_{s}(m_{h}^{2})}{4\pi}\right)^{3}A^{\textrm{PS},(3)}_{qq,h}(x)$ (65)
@@
where the 2-loop OMEs $A^{(2)}_{ij}$ have been calculated in Refs. [57, 58] while the most recent 3-loop calculation is presented in Refs. [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83].
@@ VII Results
$x$ xuvxuv xdvxdv xLxL- $xL+$ $xs+$ $xc+$ $xb+$ xgxg
@@
$x$ xuvxuv xdvxdv xLxL- $xL+$ $xs+$ $xc+$ $xb+$ xgxg
@@
In Table 2, we report the output results of Candia-v2 at aN3LO at an evolution scale $\mu_{F}$ of 100 GeV, obtained with the approximate kernels discussed above. The evolution is performed by using a logarithmic $x$-grid with 2000 grid points while each gaussian quadrature integration uses 50 gaussian points. We use $s=20$ iterations in both the NS and singlet sectors, and choose the truncation index $\kappa=30$ for the truncated solution in the singlet sector. The results exhibit negligible variation with this parameter choice. In Fig. 1, we illustrate N3LO/NNLO PDF ratios in the VFNS scheme with heavy-quark matching conditions at N3LO, while in Fig. 2 the same PDF ratios are illustrated with matching conditions at NNLO, with the $A_{ij}^{(2)}$ OMEs only, as in Refs. [107]. This allows for a separate assessment of impact from the splitting functions and OMEs. In particular, in Fig. 1 we observe an increase of approximately $6\%$ in the $c_{+}$ distribution in the $10^{-2}\lesssim x\lesssim 0.1$ range, and an increase of approximately $2.5\%$ in the $b_{+}$ combination in the $10^{-3}\lesssim x\lesssim 0.1$.
@@
In Table 3, we compare the results of Candia-v2 to those from the recent Hoppet-v2 computer code release, documented in Ref. [107]. The $x$-space Hoppet’s algorithm is based on higher-order Runge-Kutta methods for both the $x$-integrations and $Q$ evolution. Table 3 shows percentage errors, where the Hoppet-v2 result is used as the reference. The differences between Candia-v2 and Hoppet-v2 for the N3LO evolution with approximate splitting functions are overall at the sub-percent level, except for the $xc_{+}$ distribution at very large $x$, and $xb_{+}$ where differences are approximately $1\%$ in the $10^{-4}\lesssim x\lesssim 0.3$ range. The $xL_{-}$ combination at $x\lesssim 10^{-4}$ has differences that are between 1.6% and 2.3%.
@@ VIII Numerical Optimizations
Moreover, all convolutions are performed using Gauss–Legendre quadrature with a fixed set of weights and abscissae. Consequently, the splitting functions need to be evaluated only at a finite, fixed set of points—the quadrature abscissae. This allows the splitting functions evaluated at these abscissae to be cached so that, in subsequent convolutions, the cached values can be reused instead of repeatedly recomputing the splitting functions, which becomes extremely costly at higher perturbative orders.
The final major optimization exploits the fact that, unlike the coupled singlet PDFs, the non-singlet PDFs are completely decoupled. This property allows the evolution of the non-singlet PDFs to be parallelized by assigning them to separate threads. Apart from the overhead associated with thread creation and management, this strategy can in principle yield an approximately $N$-fold speedup in the non-singlet sector when using $N$ threads. After implementing these optimizations, Candia-v2 exhibits significantly improved performance relative to Candia at higher perturbative orders.

### v3

@@ I Introduction
The collinear substructure of the nucleon is described by parton distribution functions (PDFs), which map out the longitudinal momentum fractions of its inner constituent quarks and gluons in high-energy collisions. PDFs are essential components of factorization formulae in Quantum Chromodynamics (QCD), which allow for the perturbative calculation of hadronic cross sections. A precise and accurate determination of PDFs is a major endeavor [2, 3, 4] and is necessary to advance the current precision frontier of cross-section calculations at hadron colliders. In the QCD framework of collinear factorization, PDFs depend on two variables: the parton longitudinal momentum fraction $x$ and the energy scale $Q$ of the reaction. The PDF $x$-dependence is in general extracted from global QCD analyses of experimental data for a wide variety of hadronic cross sections [2, 3, 4, 5, 6, 7]. Significant efforts are currently dedicated to extracting PDFs from lattice QCD, though their precision does not yet rival that of global analyses. Recent progress in lattice-QCD PDFs is documented in [8, 9, 10, 11, 12, 13] and references therein. The energy behavior of the PDFs instead, is predicted by perturbative QCD (pQCD) and is described by the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) [14, 15, 16, 17, 18] integro-differential renormalization group equations. The latter require knowledge of the $x$-dependent PDF functional form extracted from experimental data at an initial energy scale $Q_{0}$.
@@
Parton evolution is crucial for the success of all these programs and past and recent literature contains multiple efforts devoted to the development of algorithms using different approaches to solve DGLAP equations [91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109]. In the past three decades, part of these methods have been implemented in efficient computer codes, some of which are publicly available. Examples of these codes are partonevolution [104] (Mellin-transform method with an optimized contour), Pegasus [110] ($U$-matrix method in Mellin space), Candia [111, 112] (logarithmic recursion relations in the $x$-space), Hoppet [113, 114] (Runge-Kutta method in the $x$-space), QCDNum [115] (polynomial spline interpolation in the $x$-space), and more recently Apfel [116, 117] (Runge-Kutta method and higher-order interpolations in the $x$-space), ChiliPDF [118] (global interpolation with Chebyshev polynomials in Mellin space), EKO [119] (evolution kernel operator method in Mellin space), uPDFevolv [120, 121] (parton branching method). A recent benchmarking exercise of QCD evolution at aN3LO is reported in Ref. [122].
@@ III N3LO Non-Singlet Exact Solutions in Power Series
$Z_{22}(x)$ $=\frac{1}{\beta_{3}\gamma}\left[32\pi^{2}P^{(1)}(x)+16\pi r_{1}P^{(2)}(x)-8(\bar{c}+\bar{b}r_{1})P^{(3)}(x)\right],$
@@
It may seem that any of the recursion relations in Eqns. (37a-37d) could in principle be used to determine a specific coefficient $D^{s}_{t,m,n}$. However, each recursion relation requires that particular other coefficients are known from previous steps, with the constraint that the initial condition $D^{0}_{0,0,0}$ is the only starting point for the evaluation. Therefore, there is only one procedure that can be followed for a given value of the outermost coefficient $s$ to determine all of the others:
@@ IV Truncated solutions for the DGLAP Singlet and Non-Singlet sectors at all orders in QCD
These recursion relations generate truncated solutions at arbitrary order $\kappa$ that are equivalent those obtained using the $U$-matrix method [100, 110, 105].
@@ V Factorization/Renormalization Scale Dependence
$P^{(1)}\left(x,\frac{\mu_{F}^{2}}{\mu_{R}^{2}}\right)=$ $P^{(1)}(x)-\frac{\beta_{0}}{2}P^{(0)}(x)L,$ (53)
$P^{(2)}\left(x,\frac{\mu_{F}^{2}}{\mu_{R}^{2}}\right)=$ $P^{(2)}(x)-\beta_{0}LP^{(1)}(x)+\frac{1}{4}(\beta_{0}^{2}L^{2}-\beta_{1}L)P^{(0)}(x),$ (54)
$P^{(3)}\left(x,\frac{\mu_{F}^{2}}{\mu_{R}^{2}}\right)=$ $P^{(3)}(x)-\frac{3}{2}\beta_{0}LP^{(2)}(x)+\left(\frac{3}{4}\beta_{0}^{2}L^{2}-\frac{1}{2}\beta_{1}L\right)P^{(1)}(x)$
$+\frac{1}{8}\left(-\beta_{0}^{3}L^{3}+\frac{5}{2}\beta_{0}\beta_{1}L^{2}-\beta_{2}L\right)P^{(0)}(x),$ (55)
@@ VI Heavy-Quark matching conditions
$q_{h}^{(n_{f}+1)}(x)+\bar{q}^{(n_{f}+1)}(x)=[A_{hq}^{\mathrm{S}}\otimes q^{+,(n_{f})}](x)+[A_{hg}^{\mathrm{S}}\otimes g^{(n_{f})}](x).$ (61)
Recently, the heavy-flavor asymmetry has been computed in Ref. [136] at N3LO, whose behavior is controlled by the pure-singlet OME $A_{hq}^{\mathrm{PS},\mathrm{s},(3)}$. This gives the following new relation for the heavy-flavor difference PDF:
$q_{h}^{(n_{f}+1)}(x)-\bar{q}_{h}^{(n_{f}+1)}=[A_{hq}^{\mathrm{PS},\mathrm{s}}\otimes q^{-,(n_{f})}](x).$ (62)
The OMEs have the perturbative expansions given below
@@
$A^{\textrm{PS}}_{qq,h}(x)$ $=\left(\frac{\alpha_{s}(m_{h}^{2})}{4\pi}\right)^{3}A^{\textrm{PS},(3)}_{qq,h}(x)\,$, (66)
@@
$A_{hq}^{\mathrm{PS},\mathrm{s}}(x)$ $=\left(\frac{\alpha_{s}(m_{h}^{2})}{4\pi}\right)^{3}A_{hq}^{\mathrm{PS},\mathrm{s},(3)}(x)\,$, (68)
where the 2-loop OMEs $A^{(2)}_{ij}$ have been calculated in Refs. [63, 64] while the most recent 3-loop calculations are presented in Refs. [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 136].
@@ VII Results
$x$ xuvxuv xdvxdv $_{-} xL+xL_{+}$ $xs_{+}$ $xc_{+}$ $xb_{+}$ xgxg
@@
$x$ $g$ xuxu xdxd xsxs xcxc xbxb $x\bar{u}$ $x\bar{d}$ $x\bar{s}$ $x\bar{c}$ $x\bar{b}$
@@
In Table 2, we report the output results of Candia-v2 at aN3LO at an evolution scale $\mu_{F}$ of 100 GeV, obtained with the approximate kernels discussed above. The evolution is performed by using a grid split into three parts: a region from $x=10^{-5}$ to $x=0.1$ distributed logarithmically, a region from $x=0.1$ to $x=0.9$ distributed linearly, and a region from $x=0.9$ to $x=1.0$ distributed quadratically, with points packed towards large-$x$. The regions were filled with 200, 100, and 50 points respectively. We also used $s=20$ iterations in the NS and singlet sectors, and chose a truncation index of $\kappa=20$. Lastly, the convolutions were split into the same regions as above, each using 50 Gaussian points, with appropriate mappings (logarithmic, linear, and quadratic) in the corresponding regions. The results exhibit negligible variation with this parameter choice.
In Fig. 1, we illustrate N3LO/NNLO PDF ratios in the VFNS scheme with heavy-quark matching conditions at N3LO, while in Fig. 2 the same PDF ratios are illustrated with matching conditions at NNLO, with the $A_{ij}^{(2)}$ OMEs only, as in Refs. [114]. This allows for a separate assessment of impact from the splitting functions and OMEs. In particular, in Fig. 1 we observe an increase of approximately $6\%$ in the $c_{+}$ distribution in the $10^{-2}\lesssim x\lesssim 0.1$ range, and an increase of approximately $2.5\%$ in the $b_{+}$ combination in the $10^{-3}\lesssim x\lesssim 0.1$.
@@
In Table 3, we compare the results of Candia-v2 to those from the recent Hoppet-v2 computer code release, documented in Ref. [114]. The $x$-space Hoppet’s algorithm is based on higher-order Runge-Kutta methods for both the $x$-integrations and $Q$ evolution. Table 3 shows percentage errors, where the Hoppet-v2 result is used as the reference. The differences between Candia-v2 and Hoppet-v2 for the N3LO evolution with approximate splitting functions are at the sub-percent level and below for all flavors. We show individual PDF distributions rather than $\pm$ combinations to avoid the introduction of numerical errors due to large cancellations between these PDFs at low and large $x$ as well as in the computation of the percentage errors between Candia-v2 and Hoppet-v2 in the same kinematic ranges.
@@ VIII Numerical Optimizations
Moreover, all convolutions are performed using Gauss–Legendre quadrature with a fixed set of weights and abscissae. Consequently, the splitting functions and operator matrix elements need to be evaluated only at a finite, fixed set of points—the quadrature abscissae. This allows the splitting functions evaluated at these abscissae to be cached so that, in subsequent convolutions, the cached values can be reused instead of repeatedly recomputing the splitting functions, which may become costly at higher perturbative orders. This optimization achieves a 20–25% average speedup for the exact solution by efficiently handling the exponential increase in splitting function and operator matrix element evaluations. However, improvements for the truncated solution are marginal, and the method incurs a higher memory cost that scales with grid size.
The final major optimization exploits the fact that the non-singlet PDFs are completely decoupled. This property allows the evolution of the non-singlet PDFs to be parallelized by assigning them to separate threads. Apart from the overhead associated with thread creation and management, this strategy can in principle yield an approximately $N$-fold speedup in the non-singlet sector when using $N$ threads. Further, while the singlet sector evolves a set of coupled PDFs, meaning they cannot be assigned to separate threads, we can still split up the convolution into separate pieces and perform them concurrently. This is possible due to the fact that, as mentioned previously, the convolutions require information only from the previous iteration. After implementing these optimizations, Candia-v2 exhibits significantly improved performance relative to Candia at higher perturbative orders. The initial N3LO evolution code, structured as an extension of the Candia code, exhibited execution times surpassing 20 minutes for high-precision runs. Our optimized implementation reduces this latency to less than one minute while simultaneously achieving enhanced precision.